Let R be the event that a randomly chosen athlete runs. Let W be the event that a randomly chosen athlete lifts weights. Identify the answer which expresses the following with correct notation: The probability that a randomly chosen athlete lifts weights, given that the athlete runs. Select the correct answer below: @pWR) P(RW) P(R AND W) P/W) AND P(R) Previous Page Next Page Page 10
Timothy wants to estimate the mean number of siblings for each student in his school. He records the number of siblings for each of 75 randomly selected students in the school. What is the statistic? Select the correct answer below: the specific number of siblings for each randomly selected student the 75 randomly selected students the mean number of siblings for the randomly selected students the mean number of siblings for all students in the school all the students in the school
Two students, Stephanie and Maria, want to fnd out who has the Righel GPA wnett tonlp to each of their schools. Stephanie has a GPA of 3.85, and her school has a mean GPA of 3.1 and a standard deviation of 0.4. Maria has a GPA of 3.8, and her school has a mean of 3.05 and a standard deviation of 0.2. Who has the higher GPA when compared to each their schools? Select the correct answer below: Stephanie Maria The GPAs are equal when compared to each of their schools. There is not enough information.
Question 13 (1 point) Given the following histogram for a set of data, how many values in the data set are between 9.5 and 11.5? 8 7 6 5 4 0 6.5 7.5 8.5 9.5 10.5 11.5 12.5
0 7.5 8.5 9.5 10.5 11.5 12.5 5.5 6.5 Values Select the correct answer below: 0 11 O 18 17 O 15 ● 13 Previous Page Next Page Pa
Answer #1

right choices are

(first part) P(W|R)

this is conditional probability of weight lifter given that one runs

(second part)the mean number of siblings for the randomly selected students

statistic is the sample value whereas parameter is the population value. we want to estimate population parameter with statistic using sample value.

(third part) Maria

here we use standard normal variate z=(x-mean)/sd

for stephanie z=(3.85-3.1)/0.4=1.875

for maria z=(3.8-3.05)/0.2=3,75

(fourth part) 15

frequency of 9.5-10.5 is 7 and

frequency of 10.5-11.5 is 8

so total frequency 8+7=15